# The Internal Model Principle of Time-Varying Optimization

### v1

@@ Abstract
Time-varying optimization problems are central to many engineering applications where performance metrics and system constraints evolve dynamically with time. A number of algorithms have been proposed in recent years to solve such problems; a common feature of all these methods is that they implicitly require precise knowledge of the temporal variability of the solutions in order to exactly track the optimizers. In this paper, we seek to lift these stringent assumptions. Our main result is a fundamental characterization showing that an algorithm can track an optimal trajectory if and only if it contains a model of the temporal variability of the problem. We refer to this concept to as the internal model principle of time-varying optimization. By recasting the optimization objective as a nonlinear regulation problem and using tools from center manifold theory, we provide necessary and sufficient conditions both for an optimization algorithm to achieve exact asymptotic tracking and for such an algorithm to exist. We illustrate the applicability of the approach numerically on both synthetic problems as well as practical problems in transportation.
I Introduction
@@
Time-varying optimization problems play a central role in several scientific domains, as they underpin many important contemporary engineering problems. Examples include training in Machine Learning [1, 2], dynamic signal estimation in Signal Processing [3], trajectory tracking in Robotics [4], system optimization in Industrial Control [5], and much more. Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerge as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6]. These approaches build on the classical perspective on mathematical optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices. Recently, there has been a renewed interest in the use and development of continuous-time dynamics for optimization purposes, due to the possibility of utilizing tools from dynamical systems for their analysis [7, 8] and motivated by practice, where optimization is increasingly used to control physical systems [9].
@@
Motivated by these recent developments, in this paper we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions. Although, by now, several approaches have been developed for this purpose [10, 11, 12, 13], all these techniques implicitly require full knowledge of the temporal variability of the problem [13]. Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [9, 14]), or simply because it is unrealistic to ask for a noiseless model of how the problem changes with time. Departing from this, in this paper we pose the following question: is it possible to track (exactly and asymptotically) a minimizer of a time-varying optimization problem without any knowledge of the temporal variability of the optimization? Interestingly, we prove a fundamental result showing that tracking can be achieved if and only if the temporal variability of the problem can be ‘observed’ by the algorithm, and the latter incorporates a suitably reduplicated model of such a variability. We refer to this conclusion as the internal model principle of time-varying optimization, akin to its control-theoretic counterpart [15, 16]. Our approach relies on reinterpreting the optimization algorithm design as a nonlinear, multivariable regulation problem [17], and our analysis uses tools from center manifold theory [18, 16]. Figure 1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
@@
Related works. The literature on methods for time-varying optimization is mainly divided into two classes of solutions. The first class consists of methods that do not utilize any model of the temporal variability of the problem [19, 20, 21, 22]; instead, they seek to solve a sequence of static problems. Several established approaches belong to this class, including the online gradient descent method [23] and the online Newton step algorithm [24] – see [25] and references therein. Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain ‘regret.’ Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches. In contrast, the second class of methods uses a model of the temporal evolution of the problem to seek to exactly track optimal trajectories [26, 2, 27, 28]. Particularly celebrated is the prediction-correction algorithm (see [10] and the recent work [11]), whereby at each time a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem. We refer to [13] for a recent overview of the topic. Recent years have witnessed a growing interest in this problem: [4] uses contraction to study these methods; [29] uses sampling to estimate the temporal variability of the problem; a recent survey on analyzing optimization algorithms using tools from control has appeared in [30]; constrained optimization problems are studied in [31]. Despite these efforts, the fundamental question of what is the basic algorithm structure that enables exact asymptotic tracking remains unanswered.
@@
Contributions. This paper makes four main contributions. First, we recast the problem of designing a time-varying optimization algorithm as a nonlinear multivariable regulation problem, whereby an algorithm is to be designed to regulate the gradient of a certain function to zero. We leverage this formulation to characterize the class of optimization algorithms that can achieve tracking. In a net departure from existing approaches (e.g., [23, 13, 11, 12]), our characterization is general and allows us to study not only a single optimization method, but an entire class, which enables us to derive fundamental results for all algorithms in the class. Second, by harnessing tools from center manifold theory [18, 16], we provide necessary and sufficient conditions for an optimization algorithm to ensure tracking. Interestingly, these conditions depend on the properties of the loss function (through a gradient invertibility-type condition) and on the inner model describing the temporal variability of the problem. This property allows us to prove the internal model principle of time-varying optimization, which states that for an optimization algorithm to achieve asymptotic tracking, it must incorporate a reduplicated model of the temporal variability of the problem. This feature is implicit in all existing approaches for time-varying optimization [13] but, to the best of the authors’ knowledge, lacked a formal proof until now. Third, we derive necessary and sufficient conditions for the existence of a tracking algorithm. Fourth, we use our characterizations to design algorithms for time-varying optimization. With respect to the existing literature, our algorithm does not require one to know or measure exactly the temporal variability of the problem, and thus uses less stringent assumptions. Finally, we illustrate the applicability of the approach numerically on both synthetic problems as well as practical problems in transportation.
@@
Organization. Section II presents the problem of interest, Section III studies the simpler problem of parameter-feedback optimization, which allows us to derive the necessary framework, and Section IV tackles the time-varying optimization design problem in its full generality and contains our main results. Section V discusses the tracking accuracy in relationship to the fidelity of the internal model, Section VI presents some extensions to constrained optimization, Section VII validates numerically the results, and Section VIII illustrates our conclusions. Finally, in the Appendix we summarize basic facts on center manifold theory that are used extensively throughout the paper.
@@
II Problem setting
II-A Problem statement
@@
where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:{}^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq{}^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following assumptions throughout.
@@
Assumption 2 (Existence of an exosystem).
@@
$\dot{\theta}(t)=s(\theta(t))$ (2)
for all $t\in\mathbb{R}_{\geq 0}.$square
Convexity and smoothness are standard assumptions in optimization [25]. Assumption 2 ensures the existence of an autonomous system, called exosystem, describing the class of temporal variabilities of the cost taken into consideration.
In what follows, we will say that $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow{}^{n}$ is a critical trajectory of (1) if it satisfies:
$0=\nabla_{x}f(x^{\star}(t),\theta(t)),\quad\forall t\in\mathbb{R}_{\geq 0}.$
We will assume existence of a critical trajectory and that any critical trajectory is continuous.
Remark 1.
The time variation in (1) is captured implicitly through the parameter vector $\theta(t)$. A related, yet slightly more general, problem is as follows:
@@
where the dependency on time is explicit. Problem (1) can be recast uniquely as in (3) by letting $f_{0}(x,t)=f(x,\theta(t))$ for all $x$ and $t$. On the other hand, in general, there exists an infinite number of ways to parametrize (3) as in (1), thus leading to possible ambiguities. For instance, any function $f_{0}(x,t)$ may be parametrized by $\theta(t)=t$ (so that $f_{0}\equiv f$), although this is not always compatible with our assumptions (see Assumption 3 and Remark 2 below). $\square$
In this work, we are interested in the problem of designing an optimization algorithm that is capable of determining and tracking a critical trajectory of (1). We are interested in doing so with as little knowledge as possible on $\theta(t)$ and its temporal variability. From an optimization perspective, three types of assumptions may be considered (presented in increasing order of restrictiveness):
  1. (A1)

### v2

@@ Abstract
Time-varying optimization problems are central to many engineering applications, where performance metrics and system constraints evolve dynamically with time. Several algorithms have been proposed to address these problems; a common characteristic among them is their implicit reliance on knowledge of the optimizers’ temporal variability. In this paper, we provide a fundamental characterization of this property: we show that an algorithm can track time-varying optimizers if and only if it incorporates a model of the temporal variability of the optimization problem. We refer to this concept as the internal model principle of time-varying optimization. Our analysis relies on showing that time-varying optimization problems can be recast as output regulation problems and, by using tools from center manifold theory, we establish necessary and sufficient conditions for exact asymptotic tracking. As a result, these findings enable the design of new algorithms for time-varying optimization. We demonstrate the effectiveness of the approach through numerical experiments on both synthetic problems and the dynamic traffic assignment problem from traffic control.
{IEEEkeywords}
Time-varying optimization, Internal model principle, Gradient regulation, traffic control.
1 Introduction
[…]
@@ 1 Introduction
Time-varying optimization problems play a central role in several scientific domains, as they underpin many important contemporary engineering problems. Examples include training in Machine Learning [1, 2], dynamic signal estimation in Signal Processing [3], trajectory tracking in Robotics [4], system optimization in Industrial Control [5], and much more. Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerged as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6]. These approaches build on the classical perspective on mathematical optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices. In recent years, however, numerous optimization algorithms have been formulated and studied as continuous-time processes [7, 8, 9], mainly due to the wide availability of tools for Ordinary Differential Equations (ODEs) that can facilitate their analysis [9].
@@
Motivated by these recent developments, in this paper, we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions. Several approaches have already been developed for this purpose [10, 11, 12, 13], yet all these techniques implicitly require full knowledge of the temporal variability of the problem [13]. Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [14, 15]), or simply because it is unrealistic to ask for a noiseless model of how the problem changes with time. Departing from this, in this paper we pose the following question: is it possible to track (exactly and asymptotically) a minimizer of a time-varying optimization problem without any knowledge of the temporal variability of the optimization? Interestingly, we prove a fundamental result showing that tracking can be achieved if and only if the temporal variability of the problem can be ‘observed’ by the algorithm, and the latter incorporates a suitably reduplicated model of such a variability. We refer to this conclusion as the internal model principle of time-varying optimization, akin to its control-theoretic counterpart [16, 17]. Our approach relies on reinterpreting the optimization algorithm design as a nonlinear, multivariable regulation problem [18], and our analysis uses tools from center manifold theory [19, 17]. Fig. 1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
@@
Related works. The literature on methods for time-varying optimization is mainly divided into two classes of solutions. The first class consists of methods that do not utilize any model of the temporal variability of the problem [20, 21, 22, 23]; instead, they seek to solve a sequence of static problems. Several established approaches belong to this class, including the online gradient descent method [24] and the online Newton step algorithm [25] – see [26] and references therein. Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain ‘regret.’ Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches. Moreover, suitable assumptions need to be made to provide convergence guarantees for these methods; we refer to [13, 26] and references therein for a detailed discussion. In contrast, the second class of methods uses a model of the temporal evolution of the problem to exactly track the optimal trajectories [27, 2, 28, 29]. Particularly celebrated is the prediction-correction algorithm (see [10], the recent work [11], and the discrete implementations [30]), whereby at each time a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem. We refer to [13] for a recent overview of the topic. Recent years have witnessed a growing interest in this problem: [4] uses contraction to study these methods; [31] uses sampling to estimate the temporal variability of the problem; a recent survey has appeared in [32]; constrained optimization problems are studied in [33]; recent works [34, 35, 36, 37] have modified these methods to achieve fixed-time convergence. With respect to this body of literature, which assumes precise knowledge of the temporal variability of the problem (see detailed discussion in Section 2.1), our focus here is to answer the following fundamental question: what is the least-restrictive information that is needed to track, asymptotically with zero error, the optimizer of a time-varying problem? To the best of the authors’ knowledge, a rigorous answer to this question is missing in the literature.
In this work, we recast the problem of designing optimization algorithms for time-varying optimization as a nonlinear output regulation problem. This connects our work with the literature on output regulation, which is a well-established area of research. Initial works can be traced back to the 1970s [16, 38, 18] focusing on linear systems, and later extended to nonlinear systems using both local [39, 40] and global [41] approaches for the analysis. In recent years, the field has received new attention using modern methods, see, e.g., [42, 43, 44] and the recent tutorial [45]. To the best of our knowledge, this is the first work in the literature that establishes a connection between output regulation and optimization algorithms design.
@@
Contributions. This paper features four main contributions. First, we recast the problem of designing a time-varying optimization algorithm as a nonlinear multivariable regulation problem, whereby the signal to be regulated is the gradient of the loss function. We leverage this formulation to derive necessary and sufficient conditions to achieve exact asymptotic tracking for a large class of optimization methods (described by sufficiently-smooth functions). In a net departure from existing approaches (e.g., [24, 13, 11, 12]), our characterization is general and allows us to study not only a single optimization method, but an entire class, which enables us to derive fundamental results for all algorithms in this class. Second, by harnessing tools from center manifold theory [19, 17], we provide necessary and sufficient conditions for an optimization algorithm to ensure tracking. Interestingly, these conditions depend on the properties of the loss function (through a gradient invertibility-type condition) and on the inner model describing the temporal variability of the problem. This property allows us to prove the internal model principle of time-varying optimization, which states that for an optimization algorithm to achieve asymptotic tracking, it must incorporate a reduplicated model of the temporal variability of the problem. This feature is implicit in all existing approaches for time-varying optimization [10, 11, 34, 36, 37] but, to the best of the authors’ knowledge, lacked a rigorous understanding until now. Third, we derive necessary and sufficient conditions for the existence of a tracking algorithm. Fourth, we use our characterizations to design algorithms for time-varying optimization. With respect to the existing literature, our algorithm does not require one to know or measure exactly the temporal variability of the problem and thus uses less stringent assumptions. Finally, we illustrate the applicability of the approach numerically on both synthetic problems and the traffic assignment problem in transportation.
@@
Organization. Section 2 presents the problem of interest, Section 3 the parameter-feedback problem (where the temporal variability of the problem can be explicitly measured), and Section 4 the gradient-feedback problem (where the algorithm has access only to first-order functional evaluations). Section 5 discusses the tracking accuracy in relationship to the fidelity of the internal model, Section 6 presents extensions to constrained optimization and discrete-time algorithms, Section 7 validates numerically the results, and Section 8 illustrates our conclusions. Finally, in the Appendix, we summarize basic facts on center manifold theory used in the paper.
@@
2 Problem setting
2.1 Optimization objectives
@@
where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:{}^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq{}^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following standard assumptions on the loss.
@@
Convexity and smoothness are standard assumptions in optimization [26], which have been widely used in works on related problems [8, 28, 14, 15, 10, 11, 12, 13, 24, 32, 33, 29, 31].
In (1), the parameter $\theta(t)$ is used to model the temporal variability of the problem. We will require that $\theta(t)$ belongs to a certain class of temporal variabilities, as specified next.
Assumption 2 (Class of temporal variabilities).
@@
$\dot{\theta}(t)=s(\theta(t)),$ (2)
for $t\in\mathbb{R}_{\geq 0}.$ Moreover, $\theta=0$ is an equilibrium of (2) and the trajectories of (2) are bounded.  $\square$
We stress that, a priori, we do not assume that $s(\theta)$ nor $\theta(0)$ are known (see Problem 2.2 shortly below for additional details). Assumption 2 characterizes the class of temporal variabilities taken into consideration. This assumption is mild, as it only requires that $\theta(t)$ is deterministic, sufficiently smooth (so that its derivative is some $C^{\infty}$ function of $\theta(t)$ as in (2)), and its trajectories remain bounded. In line with [39, 38], we call the autonomous system (2) the exosystem.
For simplicity of the presentation, we assume that $\Theta$ is some neighborhood of the origin of p. We put no restrictions on the size of this neighborhood (which is, e.g., allowed to be the entire space $\Theta={}^{p}$), and thus on the size of $\theta(t)$ nor on its temporal variation. Moreover, there is no restriction with asking that $\Theta$ contains the origin because, if $\theta(t)$ takes values in the neighborhood of any other point, such a point can be shifted to the origin through a time-invariant change of variables without altering the solutions of (1).
Remark 1 (Parametrization in time-varying optimization).
In (1), the time variation is captured implicitly by the parameter $\theta(t)$. A related, yet slightly more general, problem is:
@@
where the dependency on time is explicit. Problem (1) can be recast uniquely as in (3) by letting $f_{0}(x,t)=f(x,\theta(t))$ for all $x$ and $t$. On the other hand, in general, there exists an infinite number of ways to parametrize (3) as in (1), thus leading to possible ambiguities. For instance, any $f_{0}(x,t)$ may be parametrized by $\theta(t)=t$ (so that $f_{0}\equiv f$), although this is not compatible with our boundedness trajectory assumption (Assumption 2). $\square$
Remark 2 (Discrete-time implementations).
In this work, we pursue the continuous-time time-varying optimization problem (1), together with a continuous-time optimization algorithm. This is inspired by control applications (see [14] and references therein), where (1) models performance objectives associated with a physical plant to be controlled, and $\theta(t)$ models exogenous disturbances affecting the plant. If instead the parameter vector were to evolve in discrete time, we would have the alternative optimization problem
$\min_{x\in\mathbb{R}^{n}}f\left(x,\theta_{k}\right),$ (4)
where $k\in\mathbb{N}_{\geq 0}$ denotes time or iteration, and $f:\mathbb{R}^{n}\times\Theta\rightarrow\mathbb{R}$. While we could discretize $\theta(t)$ and then study discrete-time algorithms to solve this problem, physical control applications motivate studying the continuous-time problem directly. We provide additional details on discrete-time implementations of our method in Section 6.2. $\square$
In what follows, we say that $x^{\star}(t),$ with $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow{}^{n}$, is a critical trajectory of (1) if it satisfies:
$0=\nabla_{x}f(x^{\star}(t),\theta(t)),\quad\forall t\in\mathbb{R}_{\geq 0}.$ (5)
We assume the existence of a critical trajectory and that any critical trajectory is continuous. The existence of a critical trajectory can be guaranteed under standard assumptions on the optimization problem; for example, coercivity of the cost [46] (i.e., $f(x,\theta)\to\infty$ when $\|{x}\|\to\infty$), or (by Weierstrass’ theorem [46]) when the search domain can be restricted to a compact set without altering the optimizers. Continuity of the critical trajectories can also be ensured under standard assumptions: for example, by Berge’s theorem [46], by requiring that $f(x,\theta)$ is continuous in $\theta$.
In our analysis, we will use a critical point with $\theta$ at rest, which is a constant vector $x^{\star}_{\circ}\in{}^{n}$, defined implicitly as:
$0=\nabla_{x}f(x^{\star}_{\circ},0).$ (6)
We assume that $x^{\star}_{\circ}$ exists and is locally unique [^1] Existence and local uniqueness of $x^{\star}_{\circ}$ can be guaranteed, for instance, under the assumptions of the Implicit Function Theorem [47]. Namely, let $X_{\circ}\times\Theta_{\circ}$ be some neighborhood of $\left(x_{\mathrm{o}}^{\star},0\right)$, $x^{\star}_{\circ}$ exists and is locally unique when: (i) $f$ is $C^{1}$ on $X_{\circ}\times\Theta_{\circ}$, (ii) $x\mapsto f(x,\theta)$ is $C^{2}$ on $X_{\circ}$ for each $\theta\in\Theta_{\circ}$, and (iii) the $\left.\nabla_{xx}^{2}f(x,\theta)\right|_{x=x_{\circ}^{\star},\theta=0}$ is locally positive definite..
Remark 3 (Critical trajectory vs. critical point at rest).
We stress that $x^{\star}(t)$ and $x^{\star}_{\circ}$ are distinct quantities. Unlike the critical trajectory (defined in (5)), which denotes the optimal input to be tracked by the optimization algorithm, the critical point with $\theta$ at rest (defined in (6)) is a constant vector that need not be tracked and may never coincide with $x^{\star}(t)$$\square$
We conclude this discussion by reviewing two important classes of optimization methods commonly used to solve (1).
Remark 4 (Basic gradient flow algorithms).
The basic gradient flow algorithm [48, 26] for (1) reads as:
$\dot{x}(t)=-\eta\nabla_{x}f(x(t),\theta(t)),$ (7)
where $\eta>0$ models the algorithm’s step size. It is known [48] that this class of algorithms is capable of computing (asymptotically) critical points of (1) only when $\theta(t)$ is a constant signal; in all other cases, the algorithm converges to a neighborhood of a critical point and convergence is inexact. We refer the reader to Remark 11 for an insightful interpretation of this fundamental limitation. $\square$
Remark 5 (Prediction-correction algorithms).
An established family of methods to solve (1) is that of prediction-correction algorithms [10, 11, 34, 36, 37]. Tailored to our setting, the basic method of this class reads as:
[…]
where $\nabla_{xx}f(x,\theta)$ is the Hessian of $f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$ is the partial derivative of $\nabla_{x}f(x,\theta)$ with respect to $\theta$. Under strong convexity assumptions, this algorithm is known to converge exponentially, with zero asymptotic error, to a critical trajectory [10, 11]. Variations of this algorithm have also been proposed to achieve fixed-time or finite-time convergence [34]. We conclude by noting that implementing (8) requires (i) the cost to be twice continuously differentiable and strongly convex, (ii) knowledge of the maps $s(\theta)$, $\nabla_{x}f(x,\theta)$, $\nabla_{xx}f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$, and (iii) knowledge of $\theta(t).$square
2.2 Problem statement
In line with the literature [26, 10, 11, 34, 12], we focus on gradient-type algorithms for solving (1), which are algorithms that have access to first-order oracles of the cost; that is, functional evaluations of the map:
$(t,x)\mapsto\nabla_{x}f(x,\theta(t)).$ (9)
Remark 6 (Evaluating the gradient).
In the optimization literature [26], functional evaluations of (9) are generally obtained through two main approaches: (i) the algorithm has access to both the analytical form of $\nabla_{x}f(x,\theta)$ and $\theta(t)$ at each time $t$ (either because this signal is known or measurable), or (ii) an oracle returns evaluations of (9) (computed, e.g., through numerical differentiation). These alternatives will be discussed more in detail in (O1)–(O4), shortly below. $\square$
We let the optimization algorithm be described by a dynamic internal state $z(t),$ taking values on an open subset $\mathcal{Z}\subseteq{}^{n_{c}},n_{c}\in\mathbb{N}_{>0}$. The optimization algorithm generates a sequence of points $x(t)\in{}^{n}$ (called exploration signal), and has access to functional evaluations of (9) at these points (called gradient feedback signal):
$y(t)=\nabla_{x}f(x(t),\theta(t)).$ (10a)
Together with this gradient feedback signal, the optimization algorithm is described by [^2] While one could allow $G_{c}$ to also depend on $y$, so that $x(t)=G_{c}(z(t),y(t))$, we will show in Section 4 that this is unnecessary.
$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ $x(t)$ $=G_{c}(z(t)),$ (10b)
where $F_{c}:\mathcal{Z}\times{}^{n}\rightarrow{}^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow{}^{n}$ are functions to be designed. In the remainder, we refer to (10) as a dynamic gradient-feedback optimization algorithm. The architecture of the presented gradient-feedback algorithm is illustrated in Fig. 1. For simplicity of the presentation, we will require that $F_{c}(z,y)$ and $G_{c}(z)$ are such that [^3] Notice that this is without loss of generality since $F_{c}(z,y)$ and $G_{c}(z)$ are to be designed and $x_{\circ}^{\star}$ is known through (6).:
00 0 $=F_{c}(z_{\circ}^{\star},0)\qquad\text{and}\qquad x_{\circ}^{\star}=G_{c}(z_{\circ}^{\star}),$ (11)
for some locally unique $z_{\circ}^{\star}\in{}^{n_{c}}$. Together with (6), this ensures that $z_{\circ}^{\star}$ is an equilibrium of (10) when the gradient feedback signal $y(t)$ is equal to zero.
The dynamics of algorithm (10), coupled with the exosystem (2), have the form of a nonlinear autonomous system:
$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ (12a)
$y(t)$ $=\nabla_{x}f(G_{c}(z(t)),\theta(t)),$ (12b)
$\dot{\theta}(t)$ $=s(\theta(t)).$ (12c)
Definition 1 (Exact asymptotic tracking).
We say that (10) exactly asymptotically tracks a critical trajectory of (1) if there exists a neighborhood of $(z_{\circ}^{*},0)$ in $\mathcal{Z}\times\Theta$ such that, for each initial condition $(z(0),\theta(0))$ in the neighborhood, the solution of (12) satisfies $y(t)\to 0$ as $t\to\infty$. $\square$
By considering the entire class of optimization algorithms as in (10), and without requiring any assumption in addition to (9), we begin by posing the following fundamental question.
Problem 0 (Minimal knowledge for exact asymptotic tracking).
Consider the class of optimization algorithms (10). Determine the minimal necessary knowledge (beyond (9)), concerning the exosystem (2) and the optimization (1), needed to design an algorithm from this class that tracks, with zero asymptotic error, the critical trajectories of (1). $\square$
Before addressing this question, we first outline various pieces of information on the exosystem and objective that the algorithm may (or may not) be able to access. Two (non mutually-exclusive [^4] Notice that requiring both (E1) and (E2) is equivalent to asking that both $s(\theta)$ as well as $\theta(0)$ are known.) assumptions on the exosystem from the literature (see, e.g., [48, 26, 10, 11, 34, 12]) are:
  1. (E1)